# Generative Medical Event Models Improve with Scale

### v1

@@ Abstract
Realizing personalized medicine at scale calls for methods that distill insights from longitudinal patient journeys, which can be viewed as a sequence of medical events. Foundation models pretrained on large-scale medical event data represent a promising direction for scaling real-world evidence generation and generalizing to diverse downstream tasks. Using Epic Cosmos, a dataset with medical events from de-identified longitudinal health records for 16.3 billion encounters over 300 million unique patient records from 310 health systems, we introduce the Cosmos Medical Event Transformer (CoMET) models, a family of decoder-only transformer models pretrained on 118 million patients representing 115 billion discrete medical events (151 billion tokens). We present the largest scaling-law study for medical event data, establishing a methodology for pretraining and revealing power-law scaling relationships for compute, tokens, and model size. Based on this, we pretrained a series of compute-optimal models with up to 1 billion parameters. Conditioned on a patient’s real-world history, CoMET autoregressively generates the next medical event, simulating patient health timelines. We studied 78 real-world tasks, including diagnosis prediction, disease prognosis, and healthcare operations. Remarkably for a foundation model with generic pretraining and simulation-based inference, CoMET generally outperformed or matched task-specific supervised models on these tasks, without requiring task-specific fine-tuning or few-shot examples. CoMET’s predictive power consistently improves as the model and pretraining scale. Our results show that CoMET, a generative medical event foundation model, can effectively capture complex clinical dynamics, providing an extensible and generalizable framework to support clinical decision-making, streamline healthcare operations, and improve patient outcomes.
@@ 1 Introduction
Safe and effective medical care aims to deliver the right intervention to the right patient at the right time. In pursuit of this goal, patients, clinicians, and health system leaders seek consensus-driven guidelines, integrated data sources, and richer information that captures the full diversity of real-world healthcare. Optimal health outcomes require care that excels across, at minimum, four pillars: accurate diagnosis, reliable prognosis, individualized treatment planning, and optimized clinical workflow [1]. Succeeding across each of these pillars requires understanding a patient’s longitudinal medical history, addressing diagnostic and future uncertainty, incorporating patient values and goals, and adapting reasoning to temporal and clinical contexts.
@@
RWD and real-world evidence (RWE) offer a scalable path to personalized medical care. real-world data (RWD)‑driven insights already inform post-market safety surveillance, regulatory approvals, and guide therapeutic strategies for complex chronic diseases  [2, 3, 4, 5]. Today, using RWD to generate RWE at scale demands significant analytic expertise and manual curation, constraining its day-to-day impact at the point of care [2, 4]. Unlocking its full potential will require methods that can transform raw data into actionable insights at the point of care in a scalable, generalizable, and personalized way.
@@
Epic Cosmos [^1] https://cosmos.epic.com was created to address these challenges. A collaboration among health systems using Epic that is governed by a peer-elected council of participants, Cosmos aggregates de‑identified longitudinal health records for more than 300 million patients and 16.3 billion encounters as of August 2025, de-duplicating each patient’s records across health systems and combining them into a single, integrated longitudinal record. This platform unifies common clinical data—including laboratory results, diagnoses, medications, and procedures—and includes other data relevant to health, such as social drivers of health, cancer staging, genomic variants, and patient-reported outcomes, among many other data types. The de-identified data in Cosmos is intended to support patient care and accelerate scientific discovery. Insights from Cosmos are delivered to clinicians today at the point of care through features in Epic like the Cosmos Median Length of Stay, Look-Alikes, and Best Care Choices for My Patient™. Cosmos data and its downstream applications are only made available to health systems that contribute data to it. Cosmos has also been used to address a wide variety of research priorities [6] such as understanding large trends in healthcare [7, 8], investigating rare diseases [9, 10], and analyzing healthcare utilization [11, 12].
@@
Yet even at the scale of Cosmos, answering a single clinical question requires crafting custom cohort definitions, feature-engineering pipelines, and statistical analyses. To enable personalized medicine and RWE at scale for routine clinical decision-making, we need tools that can learn from the integrated patient record and can answer complex medical inquiries with flexibility, retrieving the right RWE to support clinical decision-making across a wide variety of contexts.
@@
Foundation models pretrained on real-world patient journeys have shown promise in addressing this problem, where a patient journey is formulated as a sequence of medical events. By learning latent representations of complete patient records, generative medical event models can provide patient‑specific predictions via simulated health timelines. By simulating multiple probabilistic timelines of a patient’s health, quantitative predictions can be made about the likelihood of events over specific time intervals. A single set of generated trajectories can flexibly address a wide range of clinical queries even without pre-specifying tasks, building task-specific models, or prompting a natural language model with tailored questions. Furthermore, because medical foundation models are trained specifically on medical event tokens, they can be more parameter and token-efficient in their representations. Generative medical event foundation models also offer extensive flexibility for forecasting future events, beyond just binary or quantitative prediction tasks; for example, they can predict the most likely order of events or generate a set of all events within a time frame from among hundreds of thousands of possibilities. Medical foundation models also provide a unique form of interpretability, in that clinicians and researchers can study individual generated trajectories of medical events to better validate and understand the sequence of events which the model predicts may lead up to downstream outcomes.
@@
Previous models, such as CLMBR [13], MOTOR [14], Foresight [15], ETHOS [16], and others (see Section 4.3) have illustrated the feasibility of this approach; however, their scopes are constrained by dataset breadth and depth, leaving the scalability of such approaches largely untested. Moreover, the choices of model size and compute have not been systematically studied and it’s unclear whether they are optimal and how they should scale with available data. This is particularly challenging for RWE studies at the population scale of Cosmos, as suboptimal model size and compute can be extremely costly and wasteful.
@@
To the best of our knowledge, Zhang et al. [17] conducted the first comprehensive study of the scaling laws on structured patient records. They observed power-law relationships among compute, model size, and pretraining data similar to the text domain, albeit with much higher token-to-parameter ratio, which might be attributable to the distinct characteristics of medical events. In this paper, we apply this same methodology to pretraining on Cosmos data, producing to date the largest scaling-law study on real-world patient journeys. The study in Zhang et al. [17] was limited to a de-identified dataset with only 70 thousand patients in emergency medicine. By contrast, our study is not only more than three orders of magnitude larger in terms of the number of patients, but also covers an extremely diverse range of patient populations and health conditions.
@@
We present CoMET, building on advances in generative medical event models by pretraining three decoder-only transformers with up to 1 billion parameters on Cosmos data. These models generate the next medical event—such as a diagnosis, medication order, lab result, the passage of time, or others—and these zero-shot generated sequences of medical events can be used to make clinical predictions. This work makes three contributions:
@@
    CoMET models: We describe the data transformation and training pipeline for medical event data on a dataset of 151 billion tokens from 115 billion medical events across 8.5 billion encounters.
@@
    Clinically relevant evaluations: We show that zero-shot generations with CoMET models demonstrate strong predictive performance on a wide variety of clinical tasks.
@@
By learning from the collective experience of care provided by the Cosmos community, CoMET captures patterns from data that are broad, rich, representative, and real. CoMET offers the potential for clinicians, researchers, and health systems to transform that experience into intelligence that uncovers new medical knowledge, enhances healthcare systems, and improves patient outcomes.
@@ 2 Results
2.1 CoMET Training and Inference
Figure 1: Overview of CoMET pretraining and inference. A patient journey is formulated as a sequence of medical events, and CoMET learns by predicting the next medical event. At inference time, CoMET is prompted with a patient’s medical event history and simulates potential future trajectories by autoregressively generating the next events. Predictions for any target in CoMET’s vocabulary are obtained from these simulated trajectories, enabling broad, out-of-the-box use on downstream tasks without task-specific fine-tuning or few-shot prompts.
The dataset used for training and evaluating CoMET is a filtered subset of Cosmos that comprises 115 billion medical events from 118 million unique patient records from January 2012 to April 2025 (see Section 5.1 and Table 4). We transformed each patient’s medical events into a chronological sequence, where events are each represented by compact tokens. Certain tokenization methods were inspired by ETHOS [16] and adapted for the scale and heterogeneity of Cosmos data. Section 5 provides additional information on pre-processing, sequencing, and tokenization of Cosmos data. We trained CoMET using the Qwen2 transformer architecture [18] with random initialization—i.e., without loading any pretrained Qwen2 weights (see Section 5.3 for training details). Three model sizes were trained, as detailed in Table 1. The optimal compute and training tokens used for each were determined by a scaling-law analysis, detailed in Section 2.6.
@@ 2.1 CoMET Training and Inference
CoMET-S 62M 90B 67M
CoMET-M 119M 160B 234M
CoMET-L 1B 1.7T 14B
Table 1: Trained compute-optimal CoMET models, with their size in parameters, number of training tokens, and amount of training compute measured in teraFLOPs (floating point operations).
Figure 2: Overview of CoMET evaluation performance. Each point shows the change in median evaluation scores for CoMET-S, CoMET-M, and CoMET-L relative to the best-performing task-specific supervised model in each of the major evaluation categories. For AUCROC and PR-AUC, positive values indicate that CoMET outperforms the task-specific model and negative values indicate under-performance while the opposite is true for MAE. CoMET’s performance improved with scale and generally matched or even outperformed the best task-specific supervised methods.
Figure 1 shows an overview of how CoMET inference works, and Figure 2 shows CoMET’s performance on a wide range of clinically relevant evaluations. The model is prompted with a patient’s longitudinal, tokenized record up to a desired time point. CoMET probabilistically generates $n$ simulations, which are then used to compute all predictions, including event probabilities, distributions, times-to-event, and outcome collections (see Section 5.5 for more details). For all evaluation tasks listed below, models were evaluated on data from the test set (see dataset construction methods in Section 5.1). Full tables of evaluation results are in Appendix B. Figure 2 provides a high-level snapshot of CoMET ’s performance across all evaluation categories, with detailed task-level results and descriptions presented in the subsections that follow.
To contextualize CoMET performance on these downstream tasks, we also trained three classes of supervised task-specific models (see Section 5.4 for details): linear and logistic regression, gradient-boosted decision trees (XGBoost), and supervised transformers trained from a random initialization. Each of these task-specific models was trained on its corresponding downstream task and evaluated using the same datasets and procedures as those applied to CoMET. For simplicity, figures only show the best-performing task-specific model.
2.2 CoMET models generate realistic medical event sequences
We evaluated alignment between a patient’s ground truth health records and CoMET’s generations in order to validate the plausibility of using these generations for more downstream predictive tasks.
@@ 2.2.1 Plausibility statistics
We first examined aggregate summary statistics over CoMET generations, prompting CoMET to produce 25 1-year generations for 20,000 patients. We measured CoMET’s ability to generate valid individual medical events that span multiple tokens, including diagnosis codes, medication codes, lab-result events, and department specialties in all encounter headers (see Appendix C and Table 6). Generated multi-token events were rarely invalid, and the error rate decreased as model scale increased. Furthermore, the prevalence of individual medical events for diagnoses, medications, labs, and procedures within one year as generated by CoMET all strongly agreed with their corresponding prevalence in the same patients’ 1-year future, as did the 1-year co-occurrence rate of pairs of medical events (see Appendix D and Table 6).
@@ 2.2.2 Encounter types and frequency
In addition to individual medical events, we measured how well CoMET probabilistically generates the number and types of medical encounters a given patient will experience within one year. Using the same model generations as Section 2.2.1, we found each patient’s probability distribution for the number of office visits, emergency visits, and hospital admissions that would occur in the next year. In Figure 3, calibration curves for CoMET-L show good calibration for predicting personalized healthcare needs. Across encounter types and counts, the expected calibration error (ECE[19] improved with model scale (Appendix B).
Figure 3: Calibration plots for encounter frequency. CoMET-L predicted the probability of how many encounters each patient will have within the next year, for three different encounter types (Office Visit, Emergency, and Inpatient). Each point represents a quantile group containing an equal number of patients with similar predicted probabilities. The horizontal position of each point reflects the group’s average predicted probability and the vertical position reflects the fraction of patients in that group with the specified 1-year count of encounters. Some lines do not span the full horizontal axis because few patients had those predicted probabilities. The diagonal line indicates perfect probability calibration.
@@ 2.2.3 Single-encounter generations
We next measured how well CoMET generates the full set of unique diagnosis, medication, lab, and procedure events that will occur during an encounter. For three different encounter types (office visits, emergency visits, and inpatient admissions), we chose 10,000 random encounters and prompted CoMET with the patient’s history up to and including the target encounter’s header (i.e. encounter type, department specialty, and any chief complaints). We compared the micro-averaged recall and precision of CoMET’s encounter predictions to reference values representing the recall and precision of simply filling the encounter with the patient’s past medical events over various lookback windows.
Figure 4 shows that across encounter types and medical event types, CoMET models demonstrated higher recall and precision than these lookback methods, and that with larger model sizes this performance showed consistent improvements as measured by precision-recall area under the curve (PR-AUC) (see Table 7). These precision-recall curves indicate CoMET generated medical events the patient had not previously had prior to the current encounter and filtered out most past events that were not likely to be repeated in the encounter.
Figure 4: Medical events predicted for single encounters. For office visit, emergency visit, and inpatient admissions, 10,000 random encounters of each were selected, and their medical events were compared to the medical events that CoMET predicted over 20 generations. The micro-averaged precisions and recalls are plotted over various thresholds for diagnosis, lab, medication, and procedure medical event types. In order to provide context on CoMET performance, we pooled the patient’s past events over various lookback windows and plotted the precision and recall for each. Higher area under each curve indicates better performance.
2.3 CoMET models can predict personalized future disease risk
We investigated CoMET’s ability to estimate future disease risk across multiple clinical domains and cohorts. Specifically, we categorized our prediction tasks into the following groups: disease-specific outcomes, acute-on-chronic events, and incident disease risk.

### v3

@@ Abstract
Realizing personalized medicine at scale calls for methods that distill insights from longitudinal patient journeys, which can be viewed as a sequence of medical events. Foundation models pretrained on large-scale medical event data represent a promising direction for scaling real-world evidence generation and generalizing to diverse downstream tasks. Using Epic Cosmos, a dataset with medical events from de-identified longitudinal health records for 16.3 billion encounters over 300 million unique patient records from 310 health systems, we introduce the Curiosity models, a family of decoder-only transformer models pretrained on 118 million patients representing 115 billion discrete medical events (151 billion tokens). We present the largest scaling-law study of medical event data, establishing a methodology for pretraining and revealing power-law scaling relationships for compute, tokens, and model size. Consequently, we pretrained a series of compute-optimal models with up to 1 billion parameters. Conditioned on a patient’s real-world history, Curiosity autoregressively predicts the next medical event to simulate patient health timelines. We studied 78 real-world tasks, including diagnosis prediction, disease prognosis, and healthcare operations. Remarkably for a foundation model with generic pretraining and simulation-based inference, Curiosity generally outperformed or matched task-specific supervised models on these tasks, without requiring task-specific fine-tuning or few-shot examples. Curiosity’s predictive power consistently improves as the model and pretraining scale. Our results show that Curiosity, a generative medical event foundation model, can effectively capture complex clinical dynamics, providing an extensible and generalizable framework to support clinical decision-making, streamline healthcare operations, and improve patient outcomes.
@@ 1 Introduction
Safe and effective medical care aims to deliver the right intervention to the right patient at the right time. In pursuit of this goal, patients, clinicians, and health system leaders seek consensus-driven guidelines, integrated data sources, and richer information that captures the full diversity of real-world healthcare. Optimal health outcomes require care that excels across at least four pillars: accurate diagnosis, reliable prognosis, individualized treatment planning, and optimized clinical workflow [1]. Succeeding across each of these pillars requires understanding a patient’s longitudinal medical history, addressing diagnostic and future uncertainty, incorporating patient values and goals, and adapting reasoning to temporal and clinical contexts.
@@
RWD and real-world evidence (RWE) offer a scalable path to personalized medical care. real-world data (RWD)-driven insights already inform post-market safety surveillance, support regulatory approvals, and guide therapeutic strategies for complex chronic diseases  [2, 3, 4, 5]. Today, using RWD to generate RWE at scale demands significant analytic expertise and manual curation, constraining its day-to-day impact at the point of care [2, 4]. Unlocking its full potential will require methods that can transform raw data into actionable insights at the point of care in a scalable, generalizable, and personalized way.
@@
Epic Cosmos [^1] https://cosmos.epic.com was created to address these challenges. A collaboration among health systems using Epic that is governed by a peer-elected council of participants, Cosmos aggregates de‑identified longitudinal health records for more than 300 million patients and 16.3 billion encounters as of August 2025, deduplicating each patient’s records across health systems and combining them into a single integrated longitudinal record. This platform unifies common clinical data—including laboratory results, diagnoses, medications, and procedures—and includes other data relevant to health, such as social drivers of health, cancer staging, genomic variants, and patient-reported outcomes, among many other data types. The de-identified data in Cosmos is intended to support patient care and accelerate scientific discovery. Insights from Cosmos are delivered to clinicians today at the point of care through features in Epic like the Cosmos Median Length of Stay, Look-Alikes, and Best Care Choices for My Patient™. Cosmos data and its downstream applications are only made available to health systems that contribute data to it. Cosmos has also been used to address a wide variety of research priorities [6] such as understanding large trends in healthcare [7, 8], investigating rare diseases [9, 10], and analyzing healthcare utilization [11, 12].
@@
Yet even at the scale of Cosmos, answering a single clinical question requires crafting custom cohort definitions, feature engineering pipelines, and statistical analyses. To enable personalized medicine and RWE at scale for routine clinical decision-making, we need tools that can learn from the integrated patient record and flexibly answer complex medical inquiries, retrieving the right RWE to support decisions across a wide variety of contexts.
@@
Foundation models pretrained on real-world patient journeys have shown promise in addressing this problem, where a patient journey is formulated as a sequence of medical events. By learning latent representations of complete patient records, generative medical event models can provide patient‑specific predictions through simulated health timelines. By simulating multiple probabilistic timelines of a patient’s health, quantitative predictions can be made about the likelihood of events over specific time intervals. A single set of generated trajectories can flexibly address a wide range of clinical queries even without prespecifying tasks, building task-specific models, or prompting a natural-language model with tailored questions. Furthermore, because medical foundation models are trained specifically on medical event tokens, they can be more parameter- and token-efficient in their representations. Generative medical event foundation models also offer extensive flexibility for forecasting future events, beyond binary or quantitative prediction tasks; for example, they can predict the most likely order of events or generate a set of all events within a time frame from among hundreds of thousands of possibilities. Medical foundation models also provide a unique form of interpretability, in that clinicians and researchers can study individual generated trajectories of medical events to better validate and understand the sequence of events that the model predicts may lead to downstream outcomes.
@@
Previous models, such as CLMBR [13], MOTOR [14], Foresight [15], ETHOS [16], and others (see Section 4.3) have demonstrated the feasibility of this approach; however, their scopes are constrained by dataset breadth and depth, leaving the scalability of these approaches largely untested. Moreover, the choices of model size and compute have not been systematically studied, and it is unclear whether they are optimal and how they should scale with available data. This is particularly challenging for RWE studies at the population scale of Cosmos, as suboptimal model size and compute can be extremely costly and wasteful.
@@
To the best of our knowledge, Zhang et al. [17] conducted the first comprehensive study of the scaling laws on structured patient records. They observed power-law relationships among compute, model size, and pretraining data similar to those in the text domain, albeit with a much higher optimal token-to-parameter ratio, which may be attributable to the distinct characteristics of medical events. In this paper, we apply the same methodology to pretraining on Cosmos data, producing to date the largest scaling-law study on real-world patient journeys. The study in Zhang et al. [17] was limited to a de-identified dataset with only a few hundred thousand patients in emergency medicine. By contrast, our study is not only more than three orders of magnitude larger in patient count, but also covers an extremely diverse range of patient populations and health conditions.
@@
We present Curiosity, building on advances in generative medical event models by pretraining three decoder-only transformers with up to 1 billion parameters on Cosmos data. These models generate the next medical event—such as a diagnosis, medication order, lab result, the passage of time, or others—and these zero-shot generated sequences of medical events can be used to make clinical predictions. This work makes three contributions:
@@
    Curiosity models: We describe the data transformation and training pipeline for medical event data on a dataset of 151 billion tokens derived from 115 billion medical events across 8.5 billion encounters.
@@
    Clinically relevant evaluations: We show that zero-shot generations with Curiosity models demonstrate strong predictive performance on a wide variety of clinical tasks.
@@
By learning from the collective experience of care provided by the Cosmos community, Curiosity captures patterns from data that are broad, rich, representative, and real. Curiosity offers the potential for clinicians, researchers, and health systems to transform that experience into intelligence that uncovers new medical knowledge, enhances healthcare systems, and improves patient outcomes.
@@ 2 Results
2.1 Curiosity Training and Inference
Figure 1: Overview of Curiosity pretraining and inference. A patient journey is formulated as a sequence of medical events, and Curiosity learns by predicting the next medical event. At inference time, Curiosity is prompted with a patient’s medical event history and simulates potential future trajectories by autoregressively generating the next events. Predictions for any target in Curiosity’s vocabulary are obtained from these simulated trajectories, enabling broad, out-of-the-box use on downstream tasks without task-specific fine-tuning or few-shot prompts.
The dataset used for training and evaluating Curiosity is a filtered subset of Cosmos that comprises 115 billion medical events from 118 million unique patient records spanning January 2012 to April 2025 (see Section 5.1 and Table 4). We transformed each patient’s medical events into a chronological sequence, where events are each represented by compact tokens. Certain tokenization methods were inspired by ETHOS [16] and adapted for the scale and heterogeneity of Cosmos data. Section 5 provides additional information on preprocessing, sequencing, and tokenization of Cosmos data. We trained Curiosity using the Qwen2 transformer architecture [18] with random initialization—i.e., without loading any pretrained Qwen2 weights (see Section 5.3 for training details). Three model sizes were trained, as detailed in Table 1. The optimal compute and training tokens used for each were determined by a scaling-law analysis, detailed in Section 2.6.
@@ 2.1 Curiosity Training and Inference
Curiosity-S 62M 90B 67M
Curiosity-M 119M 160B 234M
Curiosity-L 1B 1.7T 14B
Table 1: Trained compute-optimal Curiosity models, with their size in parameters, number of training tokens, and amount of training compute measured in teraFLOPs (floating-point operations).
Figure 2: Overview of Curiosity evaluation performance. Each point shows the change in median evaluation scores for Curiosity-S, Curiosity-M, and Curiosity-L relative to the best-performing task-specific supervised model in each of the major evaluation categories. For AUCROC and PR-AUC, positive values indicate that Curiosity outperforms the task-specific model and negative values indicate underperformance while the opposite is true for MAE. Curiosity’s performance improved with scale and generally matched or even outperformed the best task-specific supervised methods.
Figure 1 shows an overview of how Curiosity inference works, and Figure 2 summarizes Curiosity’s performance across a wide range of clinically relevant evaluations. The model is prompted with a patient’s longitudinal, tokenized record up to a desired time point. Curiosity probabilistically generates $n$ simulations, which are then used to compute all predictions, including event probabilities, distributions, times-to-event, and outcome collections (see Section 5.5 for more details). For all evaluation tasks listed below, models were evaluated on data from the test set (see dataset construction methods in Section 5.1). Full tables of evaluation results are in Appendix B. Figure 2 provides a high-level snapshot of Curiosity ’s performance across all evaluation categories, with detailed task-level results and descriptions presented in the subsections that follow.
To contextualize Curiosity’s performance on these downstream tasks, we also trained three classes of supervised task-specific models (see Section 5.4 for details): linear and logistic regression, gradient-boosted decision trees (XGBoost), and supervised transformers trained from random initialization. Each of these task-specific models was trained on its corresponding downstream task and evaluated using the same datasets and procedures as those used for Curiosity. For simplicity, figures only show the best-performing task-specific model.
2.2 Curiosity models generate realistic medical event sequences
We evaluated alignment between a patient’s ground truth health records and Curiosity’s generations to validate the plausibility of using these generations for more downstream predictive tasks.
@@ 2.2.1 Plausibility statistics
We first examined aggregate summary statistics over Curiosity generations, prompting Curiosity to produce 25 1-year generations for 20,000 patients. We measured Curiosity’s ability to generate valid individual medical events that span multiple tokens, including diagnosis codes, medication codes, lab-result events, and department specialties in all encounter headers (see Appendix C and Table 6). Generated multi-token events were rarely invalid, and the error rate decreased as model scale increased. Furthermore, the prevalence of individual medical events for diagnoses, medications, labs, and procedures within one year as generated by Curiosity all strongly agreed with their corresponding prevalence in the same patients’ 1-year future, as did the 1-year co-occurrence rate of pairs of medical events (see Appendix D and Table 6).
@@ 2.2.2 Encounter types and frequency
In addition to individual medical events, we measured how well Curiosity probabilistically generates the number and types of medical encounters a given patient will experience within one year. Using the same model generations as Section 2.2.1, we found each patient’s probability distribution for the number of office visits, emergency visits, and hospital admissions that would occur in the next year. In Figure 3, calibration curves for Curiosity-L show good calibration for predicting personalized healthcare needs. Across encounter types and counts, the expected calibration error (ECE[19] improved with model scale (Appendix B).
Figure 3: Calibration plots for encounter frequency. Curiosity-L predicted the probability of how many encounters each patient will have within the next year, for three encounter types (Office Visit, Emergency, and Inpatient). Each point represents a quantile group containing an equal number of patients with similar predicted probabilities. The horizontal position of each point reflects the group’s average predicted probability and the vertical position reflects the fraction of patients in that group with the specified 1-year count of encounters. Some lines do not span the full horizontal axis because few patients had those predicted probabilities. The diagonal line indicates perfect probability calibration.
@@ 2.2.3 Single-encounter generations
We next measured how well Curiosity generates the full set of unique diagnosis, medication, lab, and procedure events that will occur during an encounter. For three different encounter types (office visits, emergency visits, and inpatient admissions), we chose 10,000 random encounters and prompted Curiosity with the patient’s history up to and including the target encounter’s header (i.e. encounter type, department specialty, and any chief complaints). We compared the micro-averaged recall and precision of Curiosity’s encounter predictions to reference values representing the recall and precision of simply filling the encounter with the patient’s past medical events over various lookback windows.
Figure 4 shows that across encounter types and medical event types, Curiosity models demonstrated higher recall and precision than these lookback methods, and that with larger model sizes this performance showed consistent improvements as measured by precision-recall area under the curve (PR-AUC) (see Table 7). These precision-recall curves indicate Curiosity generated medical events the patient had not previously had prior to the current encounter and filtered out most past events that were not likely to be repeated in the encounter.
Figure 4: Medical events predicted for single encounters. For office visit, emergency visit, and inpatient admissions, 10,000 random encounters of each were selected, and their medical events were compared to the medical events that Curiosity predicted over 20 generations. The micro-averaged precisions and recalls are plotted over various thresholds for diagnosis, lab, medication, and procedure medical event types. In order to provide context on Curiosity’s performance, we pooled the patient’s past events over various lookback windows and plotted the precision and recall for each. Higher area under each curve indicates better performance.
2.3 Curiosity models can predict personalized future disease risk
We investigated Curiosity’s ability to estimate future disease risk across multiple clinical domains and cohorts. Specifically, we categorized our prediction tasks into the following groups: disease-specific outcomes, acute-on-chronic events, and incident disease risk.